OpenStudy (brittany):

what are the solutions tho this equation 6+|x+1|=12

OpenStudy (aroub):

6+x+1=12 7+x=12 x=12-7 x=5

OpenStudy (amistre64):

the absolute value notation tells us that we should get 2 results

OpenStudy (amistre64):

read it wrong ... at any rate x+1 = 6 or -6

OpenStudy (aroub):

wait.. its 5 or -5 right?

OpenStudy (anonymous):

no: 5 and -7